OpenStudy (mathhelp346):

(3r-7)+(17r-6) Explain how you got your answer too please. Thank youu!

OpenStudy (anonymous):

(3r-7)+(17r-6) 3r(17r-6)-7(17r-6) (Multiply the next step and get your answer)

jhonyy9 (jhonyy9):

write without parantheses and the same terms assum or subtract 3r-7+17r-6=3r+17r-7-6 so than this is equal 20r-13 3r+17r=20r -7-6=-13